The present age average of a couple & their daughter is 35 years.Fifteen years from now,the age of the mother will be equal to the sum of present ages of the father & the daughter . Find the present age of the mother.
(A)45 years
(B)40 years

Found 2 solutions by mananth, TimothyLamb:
Answer by mananth(16946)

let father's age be x
daughter's age be y
mother's age = z
(x+y+z)/3= 35
x+y+z=105
z= 105-(x+y)
15 years hence
105-(x+y)+15 = x+y
120-x-y=x+y
2x+2y=120
/2
x+y=60
z= 105-60
z= 45 mothers age

Answer by TimothyLamb(4379)

(m + f + d)/3 = 35
m + f + d = 105
m + 15 = f + d
---
m + f + d = 105
m + m + 15 = 105
2m = 90
---
answer:
m = 45
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php